OpenStudy (anonymous):

A Roman catapult relied on a heavy mass falling to the ground to launch a smaller mass into the air. If a mass of 200kg is allowed to fall 3m: What is the maximum energy the mass can have? If the smaller mass rose 10m what is its mass?

OpenStudy (anonymous):

This problem involves calculating the change in potential energy. The larger mass loses PE (it falls) and this energy is converted to the PE of the smaller mass. As a result, the smaller mass gains PE (it rises).

OpenStudy (anonymous):

There is a mistake in your answer. To calculate the maximum energy you have to multipy your answer with g. Without this step the measurement unit becomes wrong: J = (kg*m^2)/s^2 Your answer: kg*m Apart from that mistake a good answer.

OpenStudy (anonymous):

I made a careless mistake in the earlier answer. Here is the revised answer. PE = mgh PE (falling mass) = PE (mass being launched) m(1)*g*h(1) = m(2)*g*h(2) m(1)*h(1) = m(2)*h(2) m(1) = 200 kg and m(2) = ? h(1) = 3 m and h(2) = 10 m (200 kg) (3 m) = (m) (10 m) 600 kg*m = (m) (10 m) 60 kg = m PE = mgh = (600)(9.81) = 5886 J The max energy is 5886 J. The mass of the smaller object is 60 kg. I don't know what I was thinking earlier. Sorry! @guenny90: Thanks!